two telegraph poles "A" and "B" at the side of railway are 50m apart.a train which has a uniform acceleration passes pole "A" at speed of 70km/h,and passes "B" 2s later.what is the acceleration of the train ,and how fast is it moving as it passes pole "B"?
1
Expert's answer
2019-03-05T09:41:06-0500

(a)


"s=v_i t+\frac{at^2}{2}
\\
50=19.4\times 2+\frac{a\times 2^2}{2}
\\
a=5.6 \;\rm{m/s^2}"

(b)


"v_f=v_i+at=19.4+5.6\times 2=31 \;\rm{m/s}"